OpenStudy (anonymous):

Please help I really need help with my pre-calc homework 1. Use the rules of logarithms to expand log[(x+3)^4 (x-5)^7] 2. Use the rules of logarithms to expand ln(a^3b^5/c^8)

OpenStudy (perl):

log rules: There are three log rules you can use Log (x*y) = Log x + Log y Log (x / y ) = Log x - Log y Log (x^n) = n * Log x